OpenStudy (anonymous):

~*~*~*~HELP PLEASE ~*~*~*~ will fan and medal. please show work. Solve using quadratic formula. 6x^2+6x=-1

OpenStudy (anonymous):

\[6x ^{2}+6x=-1\]

OpenStudy (anonymous):

you should start with \[6x ^{2}+6x=-1\\6x^2+6x+1=0\] before using the formula \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]with \[a=6,b=6,c=1\]

OpenStudy (anonymous):

yes i plugged it in and got \[\frac{ -6 +- \sqrt{12} }{ 12 }\] i dont know what to do after that

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):